The activation energy for a reaction is 226.0 KJ/mol. If the rate constant is 1.20 x 10-4 S-1 at 500.0 C, what is the rate constant at 600 C
Answer #1

19号 会-完] k = 1-2 X1ct K = T. = 600 + 277 - FZZk 下, 500 +273 - 773K m| B= 226 ED : 22.6 x[of] R-F-214 丁-mail el | 226xle= * 1.